OpenStudy (anonymous):

y varies directly as the square of x. When x = 1, y = 3. What is the value of y when x = 3? 1 9 27 81

OpenStudy (anonymous):

y ~ k*x^2 3 = k*(1)^2 --> k=3 so y=3*3^2 =

OpenStudy (anonymous):

27